OpenStudy (jake124):

You decide you will buy a stock only if it shows an overall increase over the next 30 days. The first 10 days the is an average increase of $0.30. The next 10 days it has an average increase of $0.45. The next 10 days it has an average increase of $0.25. Will you buy the stock?

OpenStudy (luckycoins888):

Let's say the stock started out at price, X. So, after 10 days,the stock went up by (10days) times (30 cents a day) = X. So, after 10 days, the stock went up by (10 days) times (30 cents a day) = 3.00. So after 10 days, the stock price is X+X+3. The next 10 days, the stock decreases by (10 days) times (45 cents a day) = -4.50. So the new price of the stock after 20 days is (4.50. So the new price of the stock after 20 days is (X+3)−3)−4.50=X−1.50. So the price is below the starting price. Last 10 days the stock went up by (10days) times (25 cents a day) = X−1.50. So the price is below the starting price. Last 10 days the stock went up by (10 days) times (25 cents a day) = 2.5. So the final total is (X−X−1.50)+2.5=2.5=X+$1 So the stock price went up one dollar over the whole month.

OpenStudy (jake124):

@jabez177